QUESTION 1 (15 MARKS)

In a butter-packing factory, the quantity of butter packed in a day using a certain

machine is normally distributed. In a particular day, 12 packets of butter were taken at

random from this production line, and their masses, measured in grams, were:

9.50 9.50 11.20 10.60 9.90 11.10

10.90 9.80 10.10 10.20 10.90 11.00

By using these measurements:

a. Point out the estimator that is used to estimate the mean parameter.

[ 3 marks ]

b. Explore the 95% confidence interval for the mean mass produced by this

machine. [12 marks ]

[Total : 15 marks]


1
Expert's answer
2021-10-26T03:16:51-0400

a.


"\bar{x}=\dfrac{1}{12}(9.50+ 9.50+ 11.20+ 10.60+ 9.90+ 11.10"

"+10.90 +9.80 +10.10 +10.20 +10.90+ 11.00)"

"=\dfrac{124.7}{12}\approx10.39"


A point estimate of the mean of a population is determined by calculating the mean of a sample drawn from the population.

Point estimate of population mean


"\hat{\mu}=\mu_{\bar{x}}=10.39"

b.


"s^2=\dfrac{1}{12-1}((9.50-10.39)^2+(9.50-10.39)^2"

"+(11.20-10.39)^2+(10.60-10.39)^2"

"+(9.90-10.39)^2+(11.10-10.39)^2"

"+(10.90-10.39)^2+(9.80-10.39)^2"

"+(10.10-10.39)^2+(10.20-10.39)^2"


"+(10.90-10.39)^2+(11.00-10.39)^2"

"\approx0.399015"

"s=\sqrt{s^2}\approx0.631676"

The critical value for "\alpha = 0.05" and "df = n-1 = 11" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} = 2.200985."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(10.39-2.200985\times\dfrac{0.631676}{\sqrt{12}},"

"10.39+2.200985\times\dfrac{0.631676}{\sqrt{12}})"

"=(9.989, 10.791)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "9.989 < \mu < 10.791

," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "(9.989, 10.791)."